OpenStudy (anonymous):

solve the equation by finding square roots 5x^2=55

jhonyy9 (jhonyy9):

divide both sides by 5

jhonyy9 (jhonyy9):

than will get ?

OpenStudy (anonymous):

i got x= square root 13

jhonyy9 (jhonyy9):

come on i like make understanding all

OpenStudy (anonymous):

is that not the correct answer

jhonyy9 (jhonyy9):

13 how you have got it ?

jhonyy9 (jhonyy9):

so divide both sides by 5 than will get ?

jhonyy9 (jhonyy9):

5x^2 55 ----- = ----- divide both sides by 5 and will get 5 5 x^2 =11

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

so now make squarroot from both sides and willé get x_1,2=+/- sqrt11

OpenStudy (anonymous):

i keep getting 13 idk why

jhonyy9 (jhonyy9):

why 13 ? 13 *5=65 and not 55

OpenStudy (anonymous):

5x^2=55 x^2=55/5 x^2=11 x=sqrt11 x=+/_ 3.3

jhonyy9 (jhonyy9):

yes +/- sqrt11 and if you need to calcule at decimal value i think is right

OpenStudy (anonymous):

i am glad you understand :)

jhonyy9 (jhonyy9):

so than not is 65 ,yes ?

OpenStudy (anonymous):

i didnt need the decimal just the square root thanks

jhonyy9 (jhonyy9):

good luck bye

OpenStudy (anonymous):

thanks bye

OpenStudy (anonymous):

you are welcome